Distributive, Associative, Commutative, or Identity?
1) 7k + (-3k)
2) 5x + (3x + 7)
3) 11(12p)
4) -9y + 2y + 7
5) 2m - 3m - 8n + n
6) 2x + (x + y)

Answer by elima(1433)

1) 7k + (-3k)= Commutative property of addition; no matter what order you put your addends in the answer is always the same.
2) 5x + (3x + 7)=Associative Property of addition; The sum stays the same even when you group you addends differently.
3) 11(12p)= Distributive property
4) -9y + 2y + 7=Commutative property of addition
5) 2m - 3m - 8n + n=Commutative property of addition
6) 2x + (x + y)= Associative Property of addition